### v2

# https://www.outerlands.io/

Abstract
Optimistic rollups are a popular and promising method of increasing the throughput capacity of their underlying chain. These methods rely on economic incentives to guarantee their security. We present a model of optimistic rollups that suggests that the incentives are not necessarily aligned with the expected behavior of the players, thus potentially undermining the security of existing optimistic rollups. We discuss some potential solutions illuminated by our model. ollup, Blockchain, Scaling
Keywords:
R
1 Introduction
Blockchains are public ledgers used to record transactions and execute decentralized programs. Rollups are a ‘Layer 2’ (L2) scaling solution that operate on top of a main blockchain [16]. They add throughput capacity by moving the computations necessary for transactions off the mainchain on to the L2 chain. The data is still posted on the mainchain; a fact that constitutes the main difference between rollups and other scaling solutions. There are two major types of rollups: zero knowledge rollups and optimistic rollups. Optimistic rollups function by ‘optimistically’ expecting that the commit transactions to a blockchain is being done accurately and honestly in general [16, 5]. The execution of these commitments is overseen by other nodes, but they are not necessarily systematically checked. These differ from zero knowledge rollups, which require every commitment to be proved correct every time.
Optimistic rollups function by having two types of nodes. There are nodes that collect and perform the calculations for the bundle of transactions. They must put up a stake to participate. We will call these nodes aggregators, but note that this role is sometimes also called the sequencer. The work done by these nodes is then checked by other nodes that watch the chain and submit fraud proofs when they suspect an inaccuracy. This second role is called a validator. These validators must make a stake before they can press a challenge and they get rewarded if the challenge is successful and penalized if it is not. If a challenge is successful, the validator receives some of the aggregator’s stake and the rest is burned. If a fraud proof is brought, it must be settled on the mainchain, which can be costly, so the system is intended to disincentivize unnecessary proofs [16, 5].
Currently, the most popular implementations of optimistic rollups are Arbitrum and Optimism, with approximately 65% and 19% of the market share, respectively [12]. They differ in their precise implementation of the fault proof procedure. Arbitrum has a multi-step procedure that is designed to lessen the load on the main chain, although it still retains the same staking mechanism. Optimism has a simple procedure for resolving fault proofs, but it is more onerous on the main chain and thus more expensive. Neither implementation has decentralized aggregator, although they both state in their roadmap that decentralization is a priority and coming soon. Moving forward, we will assume that the aggregator is a decentralized and not a trusted party, but will remain agnostic on how precisely the fault proof is executed.
Optimistic rollups require users to wait to withdraw their deposits until a ‘challenge period’ has ended. This allows time for the validator nodes to check the transactions. This period is usually seven days. There are proposed solutions to this issue, for example the use of liquidity pools to allow users to access funds before their transactions are final, but the waiting period will always be an inherent part of the protocol.
1.1 Related Work
A concurrent work by Li[9] engages with a similar question. In the working paper, the author presents a model of a potential attack on the well functioning of optimistic rollups and concludes that their current design is not secure. Here, we consider a mechanism to be secure only if there is a pure equilibrium that is the intended behaviour. There are some substantial differences between their model and ours. We assume that the choice on the part of the validator of whether or not to perform the search is distinct from the choice of whether or not to challenge, i.e. a validator can blindly challenge without having performed a search. We also assume different payoffs. If a dishonest transaction is caught, we assume the aggregator does not get the utility of the attack, that is, the faulty commitment is revoked. So a caught dishonest aggregator has their stake slashed while also losing the utility of the attack. Furthermore, we assume that an erroneous challenge results in a penalty to the validator.
While Li’s model and the model we present are different, the general conclusion, that optimistic rollups are not incentive compatible, is similar. We believe that our model captures the phenomenon in a more nuanced and accurate way. The paper by Li [9] also extends their model to a case with multiple of each type of player. A similar analysis with many players is was done by researchers at Offchain Labs in another concurrent work. They also find an undesirable equilibrium when there is only one of each player [10].
Other works have used rational and economic augments to investigate other blockchain designs. This type of analysis has been applied to Byzantine Fault Tolerance [1, 7], proof-of-stake [14, 15], sharding [11], and frequently to proof-of-work systems [8, 3, 6]. A survey paper details many different applications of game theory to blockchain scaling mechanisms [16].
There has also been a great deal of research into the improvement and functioning of scaling solutions [4]. The work in Thibault et al [16] discusses many types of blockchain scaling solutions and we follow their account of the functioning of optimistic rollups in the setup of our model.
2 Our Model
Our model is set up with three players: the aggregator $A$, the validator $V$, and the transactor $T$. The transactors are the owners of the transactions that are to be committed. These are not necessarily transactions; they may instead be part of the use of a smart contract. The transactor reaps some utility $u_{T}$ from the execution of their transaction and pays fee $f$, which goes to the aggregator as compensation for their services. We assume that the utility of the transaction is greater than the fee, i.e. $u_{T}>f$. In order to commit the transaction, the aggregator must perform some underlying computational work that incurs cost, $w$. This is a computational cost, not a monetary cost, but we will assume that this and other computational costs are easily transformed into monetary costs. We will assume that the agents enjoy simple linear utility. Thus, in this stylized model of a transaction, we have the following utilities, if all goes well:
$A:$ $\quad f-w$
$V:$ 00$\quad$ 0
$T:$ $\quad u_{T}-f$
The algorithm functions as follows.
Algorithm 2.1
Protocol.
  • Transactor $T$ sends their payment $f$ and their transaction information to the aggregator $A$.
  • The aggregator stakes $s_{A}$
  • Then $A$ performs the computation of the transaction, with cost $w$ and commits the result.
  • Then $V$ monitors the integrity of the transaction:
    • If $V$ finds an error, it stakes $s_{V}$ and the proof is checked.
        If the proof is found to be correct, the payment and stake can be withdrawn by $A$ and $V$’s stake is slashed.
        If the proof is found to be incorrect, $T$ is able to withdraw their $f$ and half the aggregator stake $s_{V}$ is given to $V$ and the other half is burned.
    • If $V$ does not find an error after seven days, the aggregator is able to withdraw $f$ and $s_{A}$.
Algorithm 1: Optimistic rollup
With this algorithm in hand, we can start to look at the incentives for each party. Initially, we will assume that the validator nodes incur no cost to keep an eye on the activities on chain, but we will introduce this search cost in the next section.
One important question is what happens when something goes wrong. If a lazy aggregator acts dishonestly, they might just abscond with the fee $f$. Rather than being content with just the fee, the aggregator could siphon currency from a smart contract. This need not come from the transactor, but the ecosystem at large. A particularly large sum might be siphoned from a bridge contract, which is how the L2 connects to the L1 or another chain. One attack, though different means than those detailed here, was able to steal \$326M from a Solana bridge contract [13]. We call this potentially large benefit $z$. Attempting this attack will require about the same computational work $w$. With these assumptions, we can readily write down the normal form game and find the equilibrium.
Lemma 1
The optimistic rollup paradigm as described in Algorithm 1 yields ($P$ behaving honestly) and ($V$ posing no challenge) as the pure-strategy Nash equilibrium when $V$ has perfect information and all players are rational and risk neutral.
Proof
We have the game in extensive form as follows, with utility vectors of the form $\begin{pmatrix}A\\ V\\ T\end{pmatrix}$:
{istgame}
[…]
Game 1: Game tree corresponding to Algorithm 1
We implicitly assume that each player also has an outside option wherein they do not partake and receive a utility of 00. Using backwards induction, we readily see that if $A$ is honest, $V$’s best response is to not pose a challenge. If $A$ is not honest, $V$’s best response is to pose a challenge.
Given these strategies for $V$, $A$ can see that honest behaviour will result in a utility of $f-w$, which is positive under our assumptions, and dishonest behaviour will result in a utility of $-s_{P}$. Thus, $A$’s best strategy is to behave honestly. The transactor will receive $u_{T}-f$, which is assumed positive, and will therefore not chose the outside option. Thus ($A$ behaving honestly) and ($V$ posing no challenge) is the pure-strategy Nash equilibrium, as desired. ∎
If our assumptions were true to reality, this finding would support the incentive compatibility of optimistic rollups. We will now discuss the roles of search costs.
3 Search costs
Even without introducing search costs, it is easy to read off the fact that, given that honest behavior on the part of the prover is rational, $V$ will have an expected utility of 00. Why then would a validator participate in the protocol? We observe participation in real implementations and there are good reasons to participate other than the potential of the reward. These include the fact that the validator nodes might be run by people with an interest in the rollup protocol. Frequent users, for example, might want to keep an eye on the blockchain in order to be assured of the security of their own transactions. Running a node gives access to the shared ledger, which grants access to data that is important for decentralized applications or frequent traders. A simpler solution purposed by Arbitrum is to just pay validator nodes for their services [2].
Running a validator node is not as computationally intensive as other types of nodes in other applications. In order to model this accurately, we introduce a small search cost. If the validator incurs this search cost, they have perfect information about $A$’s honesty.
In order to model this addition, we introduce another action where the validator decides whether they intend to invest in the search cost or not. For simplicity, we consider this to be a binary choice. We also omit the transactors $T$ in this game for simplicity because they are not in a position to make a choice in the game. They could opt for an outside option, but we will consider this once we have explored the game with just $A$ and $V$.
In the following, the dotted lines represent an information sets. That is, the player does not know which of the two nodes they are really at. In context, $A$ will not know if $V$ is performing searches or not, and, if $V$ has chosen (No Search), they do not know if $A$ is honest.
In order to simplify the calculations, we normalize the honest behavior of the aggregator to 00. That is set $f-w=0$. We will also simplify the penalty that $A$ faces when they misbehave to be just the relatively large slashed stake $s_{A}$ and disregard the $w$ term.
The utility vectors are of the form $\begin{pmatrix}A\\ V\end{pmatrix}$.
[…]
Game 2: Extensive form game for optimistic rollup with search costs where the dotted lines indicate information sets.
The right tree corresponds to the game we have already discussed, except that $A$ does not know that $V$ is performing searches. We now assign three probabilities to the choices of the players: $V$ does no search with probability $b$ and then challenges with probability $g$; and $A$ is honest with probability $h$. All are between 00 and $1$.
Lemma 2
A rational, risk neutral $A$ will be indifferent to being honest if $g=1-\frac{s_{A}}{bs_{A}+bz}$.
Proof
It is easy to read off from Game 2 that if the aggregator is honest, they will receive 00 utility, regardless of $V$’s actions. Recall this is due to our normalization of $f-w$ to 00.
Suppose temporarily that $A$ knew that $V$ was not doing searches and was challenging randomly with a probability $g$. If they were then dishonest, they would then have utility of $g(-s_{A})+(1-g)z$. Let $b$ be the probability that $V$ blindly challenges, i.e. incurs no search costs. Honest behavior still results in 00 and dishonest behavior results in expected utility of
$b\cdot(g\cdot(-s_{A})+(1-g)z)+(1-b)(-s_{A}).$
This is the sum, weighted by $b$, for $V$ going in blind, selecting challenge with probability $g$, and by $(1-b)$ for $V$ performing a search and catching $A$ with certainty.
If we set this equal to 00, which is the expected utility to $A$ of behaving honestly, we get
$b\cdot(g\cdot(-s_{A})+(1-g)z)+(1-b)(-s_{A})=0$
$-bgs_{A}+bz-bgz-s_{A}+bs_{A}=0$
Solving for $g$ yields
$bz-s_{A}+bs_{A}=g\cdot(bs_{A}+bz)$
$g=\frac{bz-s_{A}+bs_{A}}{bs_{A}+bz}$
We can rearrange this slightly differently.
$g=\frac{bz+bs_{A}}{bs_{A}+bz}-\frac{s_{A}}{bs_{A}+bz}$
$g=1-\frac{s_{A}}{bs_{A}+bz}$ (1)
This is a line that traces out the values of $g$ and $b$ for which the aggregator is indifferent between being honest or dishonest.
In order for this to be a viable scenario, we must have that $10<b<$ 1 and $10<g<$ 1. This requirement will clearly be broken if Eq. 1 is less than 00. Thus, we have a viable $g$ if,
$0<1-\frac{s_{A}}{bs_{A}+bz}$
$\frac{s_{A}}{bs_{A}+bz}<1$
$\frac{s_{A}}{s_{A}+z}<b$
It is easy to see that this can be satisfied with a $b$ between 00 and $1$. Given that $z\gg s_{A}$, this bound is easily achieved in context. ∎
Lemma 3
There is a value for $h$, the probability with which the aggregator is honest, for which the validator is indifferent between search and no search.
Proof
Referring again to Game 2 and starting on the left side, where $V$ chooses no search, we can see that the validator reaps 00 utility if they do not challenge. In the challenge case, if $A$ is honest, $V$ has their stake burned for utility $-s_{V}$. In the dishonest case, they receive $\frac{1}{2}s_{A}$. Weighting this with the probabilities, we find the left hand side has expected utility
$g\cdot\left(h\cdot(-s_{V})+(1-h)\frac{1}{2}s_{A}\right).$
On the right hand side, $V$ can now make an informed decision. If $A$ is honest, $V$ will not challenge, resulting in utility $-x$ and if $A$ is dishonest, $V$ will always challenge for utility $\frac{1}{2}-x$, which represents the bounty less the search cost. Weighting these terms by the honesty probability yields the right hand side utility of
$h\cdot(-x)+(1-h)\left(\frac{1}{2}s_{A}-x\right)$
which can be simplified to
$(1-h)\left(\frac{1}{2}s_{A}\right)-x$
To get the whole expected utility, we weight both sides of the tree by the probability of no search $b$, which gives,
$bg\cdot\left(h\cdot(-s_{V})+(1-h)\frac{1}{2}s_{A}\right)+(1-b)\left((1-h)\frac{1}{2}s_{A}-x\right).$
We then set this to 00 to find the $h$ for which the validator is indifferent.
$bg\cdot\left(h\cdot(-s_{V})+(1-h)\frac{1}{2}s_{A}\right)+(1-b)\left((1-h)\frac{1}{2}s_{A}-x\right)=0$
$-bghs_{V}+bg\frac{1}{2}s_{A}-bgh\frac{1}{2}s_{A}+(1-b)\frac{1}{2}s_{A}-(1-b)h\frac{1}{2}s_{A}-(1-b)x=0$
$bg\frac{1}{2}s_{A}+(1-b)\frac{1}{2}s_{A}-(1-b)x=h\cdot\left(bgs_{V}+bg\frac{1}{2}s_{A}+(1-b)\frac{1}{2}s_{A}\right)$
$h=\frac{bg\frac{1}{2}s_{A}+(1-b)\frac{1}{2}s_{A}-(1-b)x}{bgs_{V}+bg\frac{1}{2}s_{A}+(1-b)\frac{1}{2}s_{A}}$ (2)
Which gives us a condition on $h$ as desired. ∎
We can combine Eq. 1 and Eq. 2 to find values for which both conditions hold. Since we have three variables but only two equations, we will not have a single point for our equilibrium.
Theorem 3.1
There are values of $s_{A}$, $s_{V}$, $z$, and $x$ that admit a mixed equilibrium that is not the same as that found in Lemma 2.1 if both $A$ and $V$ are rational and risk neutral.
Proof
We can plug Eq. 1 into Eq. 2 to attain a simplified closed form solution,
$h=\frac{\frac{1}{2}s_{A}z-(1-b)(zx+s_{A}x)}{\frac{1}{2}s_{A}z-s_{A}s_{V}+(z+s_{A})s_{V}b},$ (3)
after some tedious manipulations that we will omit. In context, it is reasonable to assume that the search costs are relatively small. If one stake is larger it would be that of the aggregator. The aggregator can extract a large amount of value, $z$. This yields the ordinal constraint:
$x<s_{V}\leq s_{A}\ll z$
We must also have that $10<h<$ 1 and $10<b<$ 1. We can think of the right hand side of Eq. 3 as a function in terms of $b$. In order for their to be a mixed equilibrium that is not a pure strategy, we need there to be a viable value of $h(b)$, i.e. between 00 and $1$, for some value $b$ where $10<b<$ 1. In order for that to be the case, we must have that the denominator is larger than the numerator and for both to have the same sign. Starting with the numerator of Eq. 3, we can see that it will be positive when
$\frac{1}{2}s_{A}z-zx-s_{A}x+(zx+s_{A}x)b>0$
$\left(\frac{1}{2}s_{A}-x+xb\right)z-s_{A}x+s_{A}xb>0$
We can see that if the term $\frac{1}{2}s_{A}-x+xb$ is positive, then, given that $z$ is very large, this inequality will be satisfied. It is easy to see that this term will be positive if
$\frac{1}{2}s_{A}-x+xb>0$
$\frac{1}{2}s_{A}>(1-b)x$
This is a reasonable restraint given that the stakes is considerably larger than the search cost, $x$.
Turning our attention to the denominator, it is easy to read off that the term will be positive if $\frac{1}{2}z>s_{V}$. This will be the case again given that $z$ very large.
We will now have that the whole term for $h$ is between 00 and $1$ if the numerator is less than the denominator. That is,
$\frac{1}{2}s_{A}z-(1-b)(zx+s_{A}x)<\frac{1}{2}zs_{A}-s_{A}s_{V}+(z+s_{A})s_{V}b$
which can be simplified to
$(b-1)(zx+s_{A}x)<zs_{V}b+(b-1)s_{A}s_{V}$
$bzx-zx+bs_{A}x+s_{A}x<zs_{V}b+bs_{A}s_{V}-s_{A}s_{V}$
Once again solving for $b$ yields
$s_{A}s_{V}-s_{A}x-xz<$ $b(zs_{v}+s_{A}s_{V}-xz-s_{A}x)$
$\frac{s_{A}s_{V}-s_{A}x-zx}{s_{A}s_{V}-s_{A}x-zx+zs_{V}}<$ $b$ (4)
We see that this constraint on be conforms to the requirement that $10<b<$ 1. Thus there are values for $g$, $h$, and $b$ all strictly between 00 and $1$ that admit an equilibrium. This is not the pure strategy equilibrium from Lemma 2.1 that is intended by the optimistic design. ∎
3.1 The Role of the Transactor
Lemma 4
A rational, risk neutral transactor will still participate if their potential utility $u_{T}$ is sufficiently high.
Proof
We must have that the transactor $T$ is willing to participate rather than take the outside option. Their utility when $A$ is honest is $u_{T}-f$. If $A$ is dishonest, but is caught, $T$ receives 00 utility. If $A$ is dishonest, but is not caught, which happens if $V$ is blind and does not challenge, $T$ receives $-f$ utility because they lose their fee. Combining this with the probability weights yields $h(u_{T}-f)-b(1-h)(1-g)f$. When this term is greater than 00, $T$ will still participate. We can use this to find a condition on $u_{T}$
$h\cdot(u_{T}-f)-b\cdot(1-h)(1-g)f>0$
$u_{T}>f+\frac{b\cdot(1-g)(1-h)}{h}f$
From this, we can see that as long as $u_{t}$ is somewhat larger than $f$, and as long as $h$ is not too small, the transactor will still participate. ∎
What we have shown is that strategies other than those we found in Lemma 2.1 constitute the equilibrium. We have found constraints on the model values such that if those constraints are met, the validator does no search $b$ of the time and, given no search, blindly challenges $g$ of the time. The validator is honest only $h$ of the time. The transactor will still participate despite this added uncertainty. The precise values of these depend on the values of the search cost, the stakes, and the potential benefit of getting away with dishonest behavior.
3.2 Remark
With Theorem 3.1, we have shown a result that has been described before, but we believe our model to be a more thorough articulation of the functioning of optimistic rollups. Our contribution further bolsters the understanding that the verifier’s dilemma poses a significant problem to the well-functioning of optimistic rollups. Of greater consequence, our work also illuminates the roles played by each factor in this problem. For example, if it were not for the validator stake $s_{V}$, or if the validator stake were very small, the lower bound on the probability of no search, $b$, would be $1$ or very close to it, as can be seen in Eq. 4. This would mean that the validator would rather incur the penalty and check randomly than incur the search cost and check with information. Random checking cannot necessarily be written off as a potential solution, as we will discuss in the next section. The precise value of $x$ is also important to the existence and severity of the deviation from the expected functionality. This again points towards a potential solution of alleviating this cost for validators, which we will also discuss in the next section. Furthermore, we can see that there are values for which the equilibrium we have found does not necessarily work. Further study of these values could yield a solution within the existing framework or better illuminate the potential risk of decentralization. All of these observations underscore the importance of the model we present here, which is, to the best of our knowledge, the most accurate and nuanced capture of the optimistic rollup mechanism.
4 Potential Solutions
The verifier’s dilemma has yet to become a practical problem because existing optimistic rollups have yet to decentralize their sequencers. Keeping this status quo is one potential solution, but asking how decentralization might be achieved without compromising the security of the chain is an important question. Another trivial solution would be to revert to a mechanism that requires the validators to check the integrity of the block every time, as is done with proof of stake L1s and zero knowledge (ZK) rollups. As ZK rollups get faster, this may become a more attractive option. We present here the outline of potential solutions that are more in keeping with the spirit of optimistic rollups.
4.1 Random Checks
If we were to stipulate that each block be checked, we would stray away from the spirit of optimistic rollups and lose their lightweight efficiency. Requiring intermittent, random checks could keep the optimistic efficiency without the hazard of the verifier’s dilemma. Whether such a set up would indeed be meaningfully more efficient than simply checking every time would depend on the relative value of the potential benefit from attack and the potential penalties. It would also depend on precisely how the random checks would be implemented. This would not necessitate fraud proofs, which are needed to find precisely where the chain became faulty, and instead would require proofs of just the block in question. These proofs could be done by nodes, which would again need incentives, or by smart contracts on the L1, which could have expensive execution that might fall on the users or be paid for with inflation. If we assume this could be done with a smart contract without too much additional cost to the user, we get a game where $\mathcal{C}$ is the contract ‘player’ in the place of $V$. Let us also define $p$, where $10<p<$ 1, to be the probability with which the contract checks the proof.
Lemma 5
There are values of the probability of check, $p$, $s_{A}$ and $z$ such that a rational, risk neutral aggregator will always be honest.
{istgame}
[…]
Game 3: Game tree illustrating a random check where utility is that of $A$ and the dotted line indicates an information set.
It is immediate that the expected income on the left child of the root is zero. Recall this is because we normalized it so. Since the left, honest choice on the part of $A$ is the desired outcome, we need a $p$ such that the right side yields negative income in expectation. Thus we would need,
$(1-p)z-ps_{A}<0$
$(1-p)z<ps_{A}$
Solving for $p$ yields,
$ps_{A}>z-pz$
$p(s_{A}+z)>z$
$p>\frac{z}{z+s_{A}}$
This is clearly in keeping with our requirement that $10<p<$ 1, but is also clear that, given $z>s_{A}$ and especially in cases where $z$ is much larger than $s_{A}$, $p$ will be close to $1$. Thus, there are only minor savings as compared to checking every time, but there are values for which such a setup would result in a rational, risk neutral $A$ behaving honestly, as desired. ∎
In this proof, we have assumed that there are only random checkers and no traditional validators. This is a viable solution, although it requires the approximate value of $z$ to be known and for the smart contract to work as described. A scenario with both random and traditional nodes might yield a different, potentially more efficient, result.
4.2 Random Mistakes
The motivating story for the verifier’s dilemma is a scenario in which the aggregator makes no mistakes and thus removes the incentive to ever check. This in turn leads to an opening for mistakes or malice on the part of the aggregator. Adding mistakes by design thus seems like a potential way to incentivize watchfulness on the part of the validators. This might pose a practical problem as presumably only some of the aggregator nodes, say those belonging to the foundation associated with the chain, would be making these purposeful mistakes and validators might then only check the blocks proposed by those nodes. Furthermore some entity would have to incur the cost of the stakes slashed when the mistakes are caught.
As we have shown in the proof for Eq. 2, there is an expression for $h$ for which the validator is indifferent between search and no search. Thus, we can presume that there are still values of $b$ and $g$ between 00 and $1$ that will yield a mixed equilibrium different from that desired. Modeling this precisely would require coefficients for both whether the block is being produced by a mistake-making $A$ and, if so, whether it elects to make a mistake. A non-mistake-making $A$ would have to take these terms into account. We leave the realization of such a model to future work, but it is far from immediate that reasonable values would yield an equilibrium in which the validator searches every time. It would likely be unreasonable for a mistake-making $A$ to make mistakes every time, as this would be tremendously expensive.
4.3 Easter Eggs
If the search costs, $x$, are relatively small, compensating the validators directly for their services seems like a tempting solution. However, it may not be clear whether or not the search is being performed correctly and checking this search would be yet another role with its own incentives. If it were possible to hide reward ‘Easter eggs,’ which could be found through proper searching but not a random challenge, this could act as a delivery mechanism for validator rewards. These rewards would be realized even when the aggregator behaves honestly.
Lemma 6
(Sketch) Compensating the validator with some term $e\geq x$, will result in the expected income for the validator of always searching being higher than not searching.
Let us add $e$, the expected value of finding the Easter egg, to the validator income in the case of search. Reading off from the proof of Eq. 2, we have that search yields more income if
$(1-h)\frac{1}{2}s_{A}-x+e>-ghs_{V}+g(1-h)\frac{1}{2}$
since $g<1$, it is clear to see that if we have $-x+e>-ghs_{V}$, the above term is always true. If we require that $e\geq x$, the result is immediate. ∎